# RAGAR: Retrieval Augment Personalized Image Generation Guided by Recommendation

### v1

@@ Abstract
Personalized image generation is crucial for improving the user experience, as it renders reference images into preferred ones according to user visual preferences. Although effective, existing methods face two main issues. First, existing methods treat all items in the user historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item. Disproportionately high weights for low-similarity items distort users’ visual preferences for the reference item. Second, existing methods heavily rely on consistency between generated and reference images to optimize the generation, which leads to underfitting user preferences and hinders personalization. To address these issues, we propose Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR). Our approach uses a retrieval mechanism to assign different weights to historical items according to their similarities to the reference item, thereby extracting more refined users’ visual preferences for the reference item. Then we introduce a novel rank task based on the multi-modal ranking model to optimize the personalization of the generated images instead of forcing depend on consistency. Extensive experiments and human evaluations on three real-world datasets demonstrate that RAGAR achieves significant improvements in both personalization and semantic metrics compared to five baselines.
1 Introduction
@@
Personalized image generation has been widely applied in scenarios such as advertising systems and chat software. It aims to render reference images into preferred ones based on user visual preferences. Existing image generation methods [???] often produce similar outputs for different users with similar input, failing to meet users’ need for personalization. As a personalized generation method, [?] extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories. This disproportionately amplifies the influence of irrelevant items, causing the extracted user preference to deviate from the true visual preferences on the reference item. Fig. 1(a) shows a comparative example, if a user historical sequence includes electronic products and clothes, and the reference item is ”short sleeves”, an overly simplistic approach might generate a striped T-shirt with the Apple logo—-an outcome that misses the user’s true preferences. By incorporating semantic-based retrieval, more relevant visual preferences, such as ”clothing, sport, pink”, can be captured. Hence, we propose the assumption that filtering the items semantically related to the reference item enhances user preference modeling.
Figure 1: A comparative example shows the effect of retrieval and recommendation by contrasting generated results with and without them.
@@
Another challenge lies in defining and evaluating personalization quantitatively. Current methods rely on either human evaluation or large multi-modal models [??], both of which are resource-intensive. Recent work [?] evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization. Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate generated images, shown in Fig. 1(b). These models provide an efficient and effective way to measure personalization. Moreover, by combining models and metrics from the recommendation domain, we validate in Sec. 2 that: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items.
Hence, the paper proposed the Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR), comprising three key modules: 1) Retrieval Module contains a correlation unit to calculate the semantic similarity between historical and reference items and a fusion unit to fuse the filtered visual features with dynamic weight. 2) Generation module applies a balance calibrator to combine general user preferences derived from the fine-tuned LLM and the retrieval-augmented user preferences to produce images. 3) Reflection module calculates the loss to balance semantics and personalization, guided by the discriminator.
In summary, the key contributions are as follows:
  • We are the first to emphasize the relationship between historical items and the reference item. The corresponding assumptions are supported by data analysis on real-world datasets in Sec. 2.
@@
  • We propose a novel model RAGAR to generate personalized images. We utilize the retrieval assumption realized by calculating the semantic similarity between items to enhance semantic consistency. Moreover, we design a discriminator to refine and assess the generated images, ensuring higher personalization quality.
  • Experiments on three real-world datasets demonstrate that our proposed method can achieve significant improvements in both personalization and semantic alignment compared to five competing methods.
2 Assumptions Validation
In this section, we validate our assumptions through data analysis on three real-world datasets: POG, ML-latest, and Stickers. [^1] Dataset descriptions are provided in Sec. 5.1. Details can be found in Table 1. To evaluate the effectiveness of the retrieval method, we use the multi-modal recommendation model [?] as the ranking model. We compare three types of user historical sequences: 1) ‘Ret’ contains the top-$k$ items selected using the retrieval method described in Sec. 4.3. 2) ‘Exp-Ret’ (Expanded Retrieval) contains the items ranked between $(k,2k]$ by the retrieval method. 3) ‘Random’ contains $k$ randomly selected items. To assess how well these sequences capture user preferences, we use Recall@10 and NDCG@10 as evaluation metrics.
Figure 2: Assumption validation results across three datasets.
The results shown in Fig. 2. ‘Random’ has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling. Our first assumption holds because both ‘Ret’ and ‘Exp-Ret’ achieve higher metrics, suggesting that items related to the reference items are crucial for modeling user preferences. Moreover, both metrics of ‘Ret’ are higher than those of ‘Exp-Ret’, which implies that the second assumption holds. Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher ranked items to enhance the user preferences modeling.
3 Related Work
Figure 3: The overview framework of the proposed RAGAR model.
Personalized Image Generation.
Advances in diffusion models (DMs) [?] have enabled high-quality image generation [????]. However, personalized image generation where user preferences guide the generation process, remains underexplored. Methods like [?] and [?] rely on natural language instructions, while [???] combine user interactions and instructions to create personalized advertisements or product images. Recent works bridge DMs and large language models (LLMs) [???], such as [?]. Building on this foundation, [?] incorporates user historical sequences to extract personalized features. However, [?] relies solely on consistency loss for optimization, overfitting the reference image’s features while neglecting personalized preferences. End-to-end multi-modal large language models (MLLMs) [???] show promise in personalizing images from textual instructions often overlook fine-grained preferences. Moreover, widely used evaluation metrics like FID [?], SSIM [?], LPIPS [?], and CLIP score [?] do not align well with human judgment, making them inadequate for assessing personalization. Meanwhile, evaluating personalization manually or using MLLMs [??] is inherently resource-intensive.
Recommendation Systems.
The goal of recommendation systems is to deliver personalized content that meets user preferences. These systems often rely on retrieval [??] and recommendation [???] techniques. Under the multi-modal setting, approaches like [?] integrate textual and visual features for enhanced commonsense reasoning, while [?] extends [?] with image understanding adapters. Specifically, [?] captures user-item relationships across modalities, enabling precise preference modeling for personalized recommendations.
4 Method
In this section, we formulate our task, introduce key notation, and outlines the RAGAR mainstream.
4.1 Task Formulation
We introduce the notation and task formulation for personalized image generation. For the user $u$, let the historical sequence be denoted as $\mathbf{S}_{u}=\{I_{1},I_{2},\cdots,I_{N-1}\}$, and let $I_{N}$ represent the reference item. Personalized image generation aims to produce an image that aligns with the user’s personalized preferences and preserves semantic consistency with $I_{N}$. Under the multi-modal setting, each item $I_{j}$ consists of an image and a text description, denoted as $\{img_{j},txt_{j}\}$.
4.2 RAGAR
@@
Our proposed method, RAGAR, is illustrated in Fig. 3. To reduce the impact of irrelevant items, we introduce the retrieval module that calculates semantic similarity scores between the reference item and historical items. Then, the module fuses the visual features of the items with the weight corresponding to the scores to produce retrieval-augmented preference features. Next, the generation module employs a fine-tuned LLM to derive general preference features from multi-modal historical sequences, corrected by the retrieval-augmented preference. A diffusion-based generator then generate images conditioned by corrected preference. Finally, we design a reflection module to evaluate the generated images, balancing personalization and semantics to ensure high-quality results.
4.3 Retrieval Module
The retrieval module aims to integrate visual features of historical items that are semantically relevant to the reference item, including the correlation unit and the fusion unit.
Correlation Unit.
The correlation unit transforms image information from historical items into textual features to compute semantic similarity scores with the reference item. Given the historical sequence $\mathbf{S}_{u}$ and reference item $I_{N}$ for the user $u$, we extract the textual description from the item images using a caption model (e.g., BLIP-2 [?]):
$cap_{i}=\textbf{Caption}(img_{i}),i=0,1,\cdots,N$ (1)
where $cap_{i}$ represents the caption of item $I_{i}$. Next, the textual description is processed into high-dimensional semantic features with a text encoder, such as CLIP [?]:
$\mathbf{E}^{sem}_{i}=\mathbf{Encoder}(cap_{i})$ (2)
@@ 4.3 Retrieval Module
where $\mathbf{E}^{sem}_{i}$ denotes the high-level semantic representation of each caption. Then, we calculate the semantic similarity between historical items and the reference item using cosine similarity:
$s_{i}=cosine(\mathbf{E}^{sem}_{i},\mathbf{E}^{sem}_{N}),i=0,1,\cdots,N-1$ (3)
where $s_{i}$ represents the similarity score for each historical item relative to the reference item. The similarity scores are sorted in descending order, and the top-$k$ items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant items in the user’s history.
Fusion Unit.
The fusion unit integrates preference features from the retrieval sequence by computing a weighted sum of visual features based on their semantic similarity scores, emphasizing high-similarity items while filtering out low-similarity ones. For the retrieval sequence $\mathbf{S}^{ret}_{u}$, we extract the visual features using a visual encoder:
$\mathbf{E}^{vis}_{i}=\mathbf{Encoder_{V}}(img_{i}),i=0,1,\cdots,N-1$ (4)
where $\mathbf{E}^{vis}_{i}$ denotes the high-level visual representation of each image. Then, the retrieval-augmented preference feature is computed as a weighted sum of these visual features and their corresponding similarity scores:
$\mathbf{P}^{ret}=\sum_{i=1}^{k}\mathbf{E}^{vis}_{i}\cdot s^{\prime}_{i}$ (5)
where $\mathbf{P}^{ret}$ represents the retrieval-augmented preference feature for the user $u$, and $s_{i}^{\prime}$ denotes normalized similarity scores.
4.4 Generation Module
The generation module bridges general user preferences with retrieval-augmented preferences to produce personalized images. It operates in three stages: 1) extracting general preference features using an LLM, 2) correcting the general preference features with the balance calibrator, and 3) generating personalized images using a diffusion-based generator.
Prompt Construction.
@@ 4.4 Generation Module
To capture user preferences from interactions, we transform interacted items into structured textual descriptions suitable for LLM analysis. Specifically, we adopt the approach from [?] that summarizes items with concise keywords to enhance interpretability and reduce extraneous information. We construct a prompt $p_{k}$ (detailed in supplementary material) populated with item captions and text to extract concise keywords with an LLM, denoted as $\mathbf{LLM}_{k}$:
$w_{i}=\mathbf{LLM}_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
@@
where $w_{i}$ is the keyword list for item $I_{i}$. We filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$. Next, we construct a prompt $p_{g}$ to capture the general user preference. To enhance the expressivity of the LLM, denoted as $\mathbf{LLM}_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${\left[IMG\{i\}\right]}_{i=1}^{L}$. The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ (detailed in supplementary material) to extract preferences.
General preference Extraction.
Given the prompt $p_{g}$, the output of $\mathbf{LLM}_{g}$ can be divided into two embeddings:
$[\mathbf{E}^{txt};\mathbf{E}^{img}]=\mathbf{LLM}_{g}(p_{g})$ (7)
@@
where $\mathbf{E}^{txt}$ represents text-related embedding and $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$ encodes image-related features. To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper to align $E^{img}$ with the image space following [?], a 4-layer encoder-decoder transformer with trainable queries and two linear layers. Then we concatenate the text-related embedding and the image-related embedding and obtain the multi-modal feature.
$\mathbf{E}^{m}=[\mathbf{E}^{txt};\mathbf{ModalMapper}(\mathbf{E}^{img})]$ (8)
@@
To maximize keyword utilization, we use the text encoder to generate keyword feature $\mathbf{E}^{key}$. We then employ a cross-attention layer to integrate $\mathbf{E}^{m}$ with $\mathbf{E}^{key}$:
$\mathbf{P}^{gen}=[Attn(\mathbf{E}^{m},\mathbf{E}^{key},\mathbf{E}^{key});\mathbf{E}^{key}]$ (9)
where $Attn$ denotes the cross attention layer to fuse features. In this way, we obtain the general preference feature $P^{gen}$.
Balance Calibrator.
@@
As mentioned in Sec. 4.3, the retrieval-augmented preference focuses primarily on item features associated with the reference item. To narrow the gap between the retrieval-augmented preference feature $\mathbf{P}^{ret}$ and the global preference feature $\mathbf{P}^{gen}$, we minimize the calibrator loss between them:
@@
This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences.
Image Generation.
With the corrected preference, we utilize the diffusion-based generator to generate images. During training, we sample $r$ random noises $\{\epsilon_{t}\}_{t=1}^{r}\sim\mathcal{N}(\mu,\sigma^{2}I)$ and apply reparameterization trick to enable gradient backpropagation through the sampling process:
$\mathbf{z_{t}}=\mu+\sigma\odot\epsilon_{t}$ (11)
The generator then produces $r$ images using the policy gradient update strategy discussed in Sec. 4.5. Additionally, the keyword feature $\mathbf{E}^{key}$ is used to generate a reference-like image, which serves as input to the ranking model for comparison with the personalized images during the reward computation.
$\begin{split}v^{key}&=Unet(\mathbf{E}^{key})\\ v^{gen}_{t}&=Unet(\mathbf{P}^{gen}+\mathbf{z_{t}}),t=0,1,...,r\end{split}$ (12)
where $v^{key}$ represents the image generated using the keyword features, and $v^{gen}$ represents the image generated using the general preference features.
4.5 Reflection Module
As the diffusion-based generator, propagating gradients backwards from the generated images is challenging [?]. To address this issue, we design a two-part reflection module that balances personalization and semantics in the generated images, ensuring both user alignment and semantic consistency.
Personalization Reflection
To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (RM) to evaluate images, reducing the reliance on expensive manual labeling. Given a set of images $\{v^{ref},v^{key},v^{gen}_{t}\}$, where $v^{ref}$ represents the reference image. By substituting visual features from this set into the sequence of original items, RM assigns the scores and rankings for each item:
$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (13)
where $\rho\in\{\rho^{ref},\rho^{key},\rho^{gen}_{t}\}$ is the rank score, and $rk\in\{1,2,3\}$ denotes the rank derived from the score. Intuitively, images that align better with the user’s personalization preference are expected to achieve higher scores and lower ranks. We then employ a reward function to guide the model toward generating images that reflect personalization preferences. Inspired by the policy gradient method [?], we accumulate rewards for each noise $\epsilon_{t}$:
$\mathcal{R}_{t}=\sum_{c\in\{ref,key\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (14)
where $\delta$ is a margin that encourages the model to generate higher-quality images, $\log_{\pi_{\theta}}$ computes probability density of the generated embedding under the Gaussian distribution.
@@ 4.5 Reflection Module
To enhance the semantic consistency, we minimize the distance between the mapper’s output in the generation module, $\mathbf{E}^{m}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$ using the semantic loss:
$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{m}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (16)
Combined with the calibrator loss, generated images align both semantically and visually with the reference.
Joint Reflection
We employ LoRA [?] to update a limited set of parameters in LLM, the Modal Mapper and the attention fusion components. The overall loss is defined as:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{cal}+\beta\cdot\mathcal{L}_{\mathrm{rank}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (17)
where $\alpha$, $\beta$ and $\gamma$ are adjustable weighting hyper-parameters.
Features POG ML-latest SER30K
#Users 1000 4689 2230
#Items 19242 9742 30739
Table 1: The characteristics of experimental datasets.
5 Experimentation
We evaluate RAGAR on three different scenarios: commodity, movie poster and sticker. Our experiments aim to answer the following research questions:
@@ 5 Experimentation
  • RQ 2: Dose RAGAR outperform baseline methods in human evaluation?
  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute in the generation?
@@
5.1 Experimental Settings
Datasets.
We utilize three real-world datasets to generate personalized images across different scenarios: POG [^2] https://github.com/wenyuer/POG.,ML-latest [^3] https://grouplens.org/datasets/movielens.,SER30K [^4] https://github.com/nku-shengzheliu/SER30K.. POG is a multi-modal dataset of fashion clothing with user interaction history. Due to the large size, we randomly select a subset of users and interactions. ML-latest is a benchmark movie dataset with user ratings. We collect corresponding posters from IMDB [^5] https://www.imdb.com. and split user historical sequences to ensure each contains 20 interactions. SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label. We merge two random themes into user historical sequences and randomly select a reference sticker from other themes. Dataset statistics are summarized in Table 1.
@@ 5.1 Experimental Settings
We compare RAGAR with five generative baselines, including three DM-based models Glide [?], SD [?] and TI [?], and two LLM-based models LaVIT [?] and PMG [?].

### v2

# RAGAR: Retrieval Augmented Personalized Image Generation Guided by Recommendation

@@ Abstract
Personalized image generation is crucial for improving the user experience, as it renders reference images into preferred ones according to user visual preferences. Although effective, existing methods face two main issues. First, existing methods treat all items in the user’s historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item. Disproportionately high weights for low-similarity items distort user visual preferences for the reference item. Second, existing methods heavily rely on consistency between generated and reference images to optimize generation, which leads to underfitting user preferences and hinders personalization. To address these issues, we propose Retrieval Augmented Personalized Image GenerAtion guided by Recommendation (RAGAR). Our approach uses a retrieval mechanism to assign different weights to historical items according to their similarities to the reference item, thereby extracting more refined users’ visual preferences for the reference item. Then we introduce a novel rank task based on the multi-modal ranking model to optimize the personalization of the generated images instead of forcing depend on consistency. Extensive experiments and human evaluations on three real-world datasets demonstrate that RAGAR achieves significant improvements in both personalization and semantic metrics compared to five baselines.
Introduction
@@
Personalized image generation has been widely applied in scenarios such as advertising systems and chat software. It aims to render reference images into preferred ones based on the user’s visual preferences. Existing image generation methods (Nichol et al. 2022; Rombach et al. 2021; Gal et al. 2023) often produce similar outputs for different users with similar input, failing to meet users’ needs for personalization. As a personalized generation method, PMG (Shen et al. 2024) extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories. This disproportionately amplifies the influence of irrelevant items, causing the extracted user preference to deviate from the true visual preferences on the reference item. Fig. 1(a) shows a comparative example, if a user historical sequence includes electronic products and clothes, and the reference item is ”short sleeves”, an overly simplistic approach might generate a striped T-shirt with the Apple logo—-an outcome that misses the user’s true preferences. By incorporating semantic-based retrieval, more relevant visual preferences, such as ”clothing, sport, pink”, can be captured. Hence, we propose the assumption that filtering the items semantically related to the reference item enhances user preference modeling.
Figure 1: Effect of retrieval and recommendation on image quality. Lower rank and higher score of RAGAR-generated images suggest a strong positive correlation between retrieval/recommendation and personalized generation quality.
Figure 2: Comparison of user history sequence construction strategies on POG. Top-$k$ retrieved items (Ret) consistently outperform lower-ranked (Exp-Ret) and random (Random) ones, validating that high-ranking relevant items better capture user preferences.
@@
Another challenge lies in defining and evaluating personalization quantitatively. Current methods rely on either human evaluation or large multi-modal models (Cho et al. 2024; Wu et al. 2024), both of which are resource-intensive. Recent work (Shen et al. 2024) evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization. Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate images, shown in Fig. 1(b). These models provide an efficient and effective way to measure personalization. Moreover, by combining models and metrics from the recommendation domain, we propose and validate two assumptions: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items. We compare three types of user historical sequences on POG (Chen et al. 2019) dataset: Ret contains the top-$k$ items selected with the retrieval method described in Sec. Retrieval Module. Exp-Ret (Expanded Retrieval) contains items ranked between $(k,2k]$ by the retrieval method. Random contains $k$ randomly selected items. The results are shown in Fig. 2. The Random sequence has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling. Our first assumption holds because both Ret and Exp-Ret achieve higher metrics, suggesting that items related to the reference items are crucial for modeling preferences. Moreover, both metrics of Ret are higher than those of Exp-Ret, which implies the second assumption holds. Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher-ranked items to enhance the preferences modeling. More results can be found in the Appendix.
Hence, we propose RAGAR (Retrieval Augmented Personalized Image Generation guided by Recommendation), comprising three key modules. The Retrieval module identifies semantically relevant historical items via a correlation unit and aggregates their visual features through a fusion unit to construct a retrieval-augmented preference. The Generation module extracts item keywords from the user sequence and integrates both detailed and global preferences, calibrated by the balance calibrator, to generate personalized images. Finally, the Reflection module applies multiple assessment signals to guide the model toward achieving a better balance between personalization and semantic fidelity.
The main contribution of the paper can be summarized as follows:
  • To our knowledge, we are the first to emphasise and explore the intrinsic relationship between historical and reference items. We hypothesise this relationship is crucial in capturing preferences more accurately and guiding the personalized generation process. To support this hypothesis, we conduct a comprehensive data analysis on real-world datasets, which reveals significant impact of item-to-item relevance on personalization quality.
@@
  • We propose a novel personalized image generation framework, RAGAR, which is specifically designed to address the dual objectives of semantic alignment and personalization. We utilize the retrieval assumption realized by calculating the semantic similarity between items to enhance semantic consistency. Additionally, we incorporate a discriminator to assess and enhance image quality, promoting better personalization.
  • Experiments on three diverse and real-world datasets demonstrate that RAGAR outperforms five strong baseline methods across multiple evaluation metrics. The empirical results highlight the effectiveness of our retrieval-based design and the discriminator.
Methods
Task Formulation
We formulate the task of personalized image generation in a multi-modal setting. Let $u$ denote a user, and let their historical interaction sequence be represented as: $\mathbf{S}_{u}=\{I_{1},I_{2},\dots,I_{N-1}\}$, where each item $I_{j}=\{\text{img}_{j},\text{txt}_{j}\}$ is a multi-modal pair consisting of an image $\text{img}_{j}$ and a corresponding textual description $\text{txt}_{j}$. The final item $I_{N}={\text{img}_{N},\text{txt}_{N}}$ serves as the reference item, which specifies the semantic content for generation. The goal is to generate a personalized image $\hat{\text{img}}_{u}$ that reflects the user’s historical preferences captured in $\mathbf{S}_{u}$ and aligns semantically with the reference item $I_{N}$. Formally, the task is to learn a generation function:
$\hat{\text{img}}_{u}=\mathcal{G}(\mathbf{S}_{u},I_{N})$ (1)
such that $\hat{\text{img}}_{u}$ maximizes personalization fidelity with respect to $\mathbf{S}_{u}$ and semantic consistency with respect to $I_{N}$.
RAGAR
@@ Methods
Our proposed method, RAGAR, is illustrated in Fig. 3. To mitigate the influence of irrelevant historical items, the retrieval module computes semantic similarity scores between the reference item and historical items by correlation unit. The fusion unit then fuses the visual features of the most relevant items, weighted by similarity scores, to construct a retrieval-augmented preference. The generation module employs a LLM to extract and summarise item keywords from the sequence. The LLM and CLIP are then used to derive detailed and global user preferences, respectively, which are subsequently integrated via a balance calibrator. A diffusion-based generator then synthesizes personalized images based on the general preference. Finally, the reflection module evaluates the generated images and updates the model by balancing semantic consistency and personalization through multi-dimensional assessments, thereby ensuring high-quality results.
Figure 3: Overview of the RAGAR framework. The framework consists of three stages. In the Retrieval stage, semantic and visual features are extracted to identify relevant historical items, whose visual features are fused into the retrieval-augmented preference. The Generation stage uses a LLM, CLIP and the balance calibrator to extract and integrate detailed and global preferences, enabling personalized image generation. In the Reflection stage, the generated images are evaluated through ranking, calibration, and semantic alignment, with policy gradients guiding the model to balance personalization and semantic fidelity.
Retrieval Module
Given the user’s interaction history $\mathbf{S}_{u}$ and a reference item $I_{N}$, we compute the semantic relevance of each historical item. First, we transform items’ image into the caption ${cap}_{i}$ using a pretrained image-to-text model (e.g., BLIP-2 (Li et al. 2023)). Each caption is then encoded into a high-dimensional semantic embedding via a text encoder such as CLIP (Radford et al. 2021):
$\mathbf{E}^{sem}_{i}=\mathbf{Enc}^{sem}(cap_{i}),\quad i=0,1,\cdots,N$ (2)
@@
We compute the semantic similarity between each historical item and the reference item using cosine similarity:
$s_{i}=\frac{\sum\mathbf{E}^{sem}_{i}\mathbf{E}^{sem}_{N}}{\sqrt{\sum{\mathbf{E}^{sem}_{i}}^{2}}\,\sqrt{\sum{\mathbf{E}^{sem}_{N}}^{2}}},\quad i=0,1,\cdots,N-1$ (3)
The top-$k$ most similar items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant preference signals from the user history. To form a retrieval-augmented preference, we extract visual features of the selected items using a visual encoder (e.g., ViT (Dosovitskiy et al. 2021)):
$\mathbf{E}^{vis}_{i}=\mathbf{Enc}^{vis}(img_{i}),\quad i=0,1,\cdots,N-1$ (4)
These top-k features are then aggregated using the normalized similarity scores $s^{\prime}_{i}$ as weights. The resulting vector $\mathbf{P}^{ret}$ captures a preference representation that emphasizes semantically aligned visual cues while suppressing irrelevant information.
$\mathbf{P}^{ret}=\sum_{i=1}^{k}\frac{\exp(s^{\prime}_{i})}{\sum_{j=1}^{k}\exp(s^{\prime}_{j})}\cdot\mathbf{E}^{vis}_{i}$ (5)
Generation Module
@@
The generation module bridges global preference with detailed preference to produce personalized images. To capture user preferences from interactions, we transform interacted items into structured textual descriptions suitable for LLM analysis. Specifically, we summary items with concise keywords to enhance interpretability and reduce extraneous information. We construct a prompt $p_{k}$ (all prompts are detailed in Appendix) populated with item text to extract concise keywords with an LLM (e.g., Qwen2.5 (Bai et al. 2023)), denoted as $\phi_{k}$:
$w_{i}=\phi_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
@@
With the keyword list $w_{i}$ for item $I_{i}$, we filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$. Next, we construct a prompt $p_{g}$ to capture the general user preference. To enhance the expressivity of the LLM, denoted as $\phi_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${[\mathbf{IMG}\{i\}]}_{i=1}^{L}$. The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ to extract preferences. Given the prompt $p_{g}$, we divide the output into text-related embedding $\mathbf{E}^{txt}$ and $\mathbf{[IMG]}$-related features $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$:
$\mathbf{E}^{txt};\mathbf{E}^{img}]=\phi_{g}(p_{g})$ (7)
@@
To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper, a 4-layer encoder-decoder transformer with trainable queries and two linear layers, to align $E^{img}$ with the image space following Gill (Koh, Fried, and Salakhutdinov 2023). Then we concatenate the text-related embedding and the image-related embedding and obtain the detailed multi-modal feature.
$\mathbf{E}^{d}=[\mathbf{E}^{txt};\phi_{m}(\mathbf{E}^{img})]$ (8)
@@
To maximize keyword utilization and capture the global features, we use the text encoder to generate keyword feature $\mathbf{E}^{g}$. We employ a cross-attention layer and a residual connection to integrate $\mathbf{E}^{d}$ with $\mathbf{E}^{g}$. In this way, we obtain the general preference feature $P^{gen}$. The generator is then used to generate personalized image.
$\mathbf{P}^{gen}=[\mathrm{softmax}(\frac{\mathbf{E}^{d}{\mathbf{E}^{g}}^{\top}}{{\sqrt{d_{\mathbf{E}^{g}}}}})\mathbf{E}^{g};\mathbf{E}^{g}]$ (9)
Reflection Module
As the diffusion-based generator, propagating gradients backwards from the generated images is challenging (Shen et al. 2024). To address this issue, we design a multi-assessment reflection module based on policy gradient (Sutton et al. 1999), ensuring both user alignment and semantic consistency.
Recommendation Reflection
To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (e.g., MICRO (Zhang et al. 2023)) to evaluate images, reducing the reliance on expensive manual labeling. More details about the ranking model (RM) can be found in Appendix. Given a set of images $\{v^{ref},v^{glob},v^{gen}\}$, where $v^{ref}$ represents the reference image, $v^{glob}$ represents the image generated with the global preference and $v^{gen}$ represents the image generated by RAGAR. By substituting visual features from this set into the sequence of original items, the RM assigns the scores $\rho\in\{\rho^{ref},\rho^{glob},\rho^{gen}\}$ and rankings $rk\in\{1,2,3\}$ derived from the score for each item:
$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (10)
Intuitively, images that better align with the user’s personalization preference are expected to achieve higher scores and lower ranks.
To enable gradient propagation from image-level feedback, we adopt a policy gradient-inspired approach. The underlying formulation is based on a Markov Decision Process (MDP), defined as $\langle\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\rangle$ (state/action spaces, reward, transition probability, discount factor). In RAGAR, we simplify the multi-step decision process into a single-step generation by setting $\mathcal{P}=\gamma=1.0$, while retaining the core policy gradient framework. Model parameters are treated as states, image generation conditioned on user preference as actions, and rank-based rewards as action values. Although the states do not transition explicitly, parameter updates implicitly induce state changes. To support multi-action exploration, we sample $r$ perturbations $\epsilon_{t}$, $t\in[1,r]$, from a Gaussian distribution over user preferences and compute the final reward as the mean rank score of the generated images. Since policy gradients require maximizing reward while other losses are minimized, we incorporate the reward as a penalty term by negating it.
Specifically, we employ a reward function to guide the model toward generating images that reflect personalization preferences. We then accumulate rewards for each noise $\epsilon_{t}$ and calculate the $\mathcal{L}_{rank}$.
$\mathcal{R}_{t}=\sum_{c\in\{ref,glob\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (11)
To encourage the model to generate higher-quality images we add a margin $\delta$. And $\log_{\pi_{\theta}}$ computes the probability density of the generated embedding under the Gaussian distribution.
Calibrator Reflection
@@
As mentioned in Sec. Retrieval Module, the retrieval-augmented preference focuses primarily on item features associated with the reference item. To narrow the gap between the general preference feature $\mathbf{P}^{gen}$ and the retrieval-augmented preference feature $\mathbf{P}^{ret}$, we minimize the calibrator loss between them:
@@
This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences, and guide the model to learn from retrieval.
@@
To enhance the semantic consistency, we minimize the distance between the modal mapper’s output in the generation module, $\mathbf{E}^{d}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$, using the semantic loss. Combined with the semantic loss, generated images align more semantically with the reference.
$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{d}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (14)
Reflection Loss
We update the Balance Calibrator in the generation module. The overall loss is defined with three adjustable weighting hyper-parameters $\alpha$, $\beta$ and $\gamma$:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{rank}+\beta\cdot\mathcal{L}_{\mathrm{cal}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (15)
Experiment
We evaluate RAGAR on three different scenarios: commodity, movie poster, and sticker. Our experiments aim to answer the following research questions:
@@
  • RQ 2: Does RAGAR outperform baseline methods in human evaluation?
  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute to the generation?
@@
Experimental Settings
Features POG ML-latest SER30K
#Users 1000 4689 2230
#Items 19242 9742 30739
Table 2: Characteristics of experimental datasets.
Datasets.
We utilize three real-world datasets to generate personalized images across different scenarios: POG, ML-latest ,SER30K. POG is a multi-modal dataset of fashion clothing with user interaction history. Due to the large size, we randomly select a subset of users and interactions. ML-latest is a benchmark movie dataset with user ratings. We collect corresponding posters from IMDB and split user historical sequences to ensure each contains 20 interactions. SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label. We merge two random themes into user historical sequences and randomly select a reference sticker from other themes. Dataset statistics are summarized in Tab. 2.
@@
We compare RAGAR with five generative baselines, including three DM-based models Glide (Nichol et al. 2022), SD (Rombach et al. 2021), and TI (Gal et al. 2023), and two LLM-based models LaVIT (Jin et al. 2024) and PMG (Shen et al. 2024).